The larger of the two numbers is 1 more than 3 times the smaller. The differences between 8 times the smaller and 2 times is 10. Find the numbers.
Answer by ankor@dixie-net.com(12701)

write an equation for each statement; Two numbers; x and y, with x being the larger
:
The larger of the two numbers is 1 more than 3 times the smaller.
x = 3y + 1
:
The differences between 8 times the smaller and 2 times is 10.
8y - 2x = 10
:
Find the numbers.
:
Replace x with (3y+1) in the above equation
8y - 2(3y+1) = 10
8y - 6y - 2 = 10
2y = 10 + 2
2y = 12
y = 6
then
x = 3(6) + 1
x = 19
;
;
Check solution in the statement:
" The differences between 8 times the smaller and 2 times is 10. "
8(6) - 2(19) = 10